Probability

Get insights from 304 questions on Probability, answered by students, alumni, and experts. You may also ask and answer any question you like about Probability

  • 304

    Questions

  • 0

    Discussions

  • 0

    Active Users

  • 4

    Followers

Have a question about Probability? Ask our experts

Ask Your Question

Keep it short & simple. Type complete word. Avoid abusive language.

New answer posted•a year ago
New answer posted•a year ago
New answer posted•a year ago
S
a year ago

This is a Long Answer Type Questions as classified in NCERT Exemplar

Sol:

Given  that  n  coins  are  two  headed  coins  and  the  remaining  (n+1)  coins  are  fair.Let  E1:the  event  that  unfair  coin  is  selected.         E2:the  event  that  fair  coin  is  selected.         E3:the  event  that  the  toss  result  is  a  head.∴     P(E1)=n2n+1  and  P(E2)=n+12n+1 P(E/E1)=1    (sure  event)  and  P(E/E2)=12∴      P(E)=P(E1).P(E/E1)+P(E2).P(E/E2)                      =n2n+1.1+n+12n+1.12=12n+1(n+n+12)                      =12n+1(2n+n+12)=3n+12(2n+1)But P(E)=3142(given)∴    3n+12(2n+1)=3142       ⇒3n+12n+1=3121                                            ⇒63n+21=62n+31                                            ⇒                n=10Hence,  the  required  value  of  n  is  10.

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a Long Answer Type Questions as classified in NCERT Exemplar

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
New answer posted•a year ago
S
a year ago

This is a Long Answer Type Questions as classified in NCERT Exemplar

Sol:

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a Long Answer Type Questions as classified in NCERT Exemplar

Sol:

Let  E1:the  event  that  the  item  is  manufactured  on  machine  A         E2:the  event  that  the  item  is  manufactured  on  machine  B         E3:the  event  that  the  item  is  manufactured  on  machine  CLet  H  be  the  event  that  the  selected  item  is  defective.  Using Bayes'  Theorm∴     P(E1)=50100,  P(E2)=30100  and  P(E3)=20100 P(H/E1)=2100,  P(H/E2)=2100  and  P(H/E3)=3100∴      P(E1/H)=P(E1).P(H/E1)P(E1).P(H/E1)+P(E2).P(H/E2)+P(E3).P(H/E3)                               =50100×210050100×2100+30100×2100+20100×3100=100100+60+60=100220=1022=511Hence,  the  required  probability  is  511.

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a Long Answer Type Questions as classified in NCERT Exemplar

Sol:

Let  E1:the  event  that  a  person  has  TB         E2:the  event  that  a  person  does  not  have  TBLet  H  be  the  event  that  the  person  is  diagnosed  to  have  TB.  Bayes'  Theorm∴     P(E1)=11000=0.001,  P(E2)=1−11000=9991000=0.999 P(H/E1)=0.99,  P(H/E2)=0.001∴      P(E1/H)=P(E1).P(H/E1)P(E1).P(H/E1)+P(E2).P(H/E2)                               =0.001×0.990.001×0.99+0.999×0.001=0.990.99+0.999                               =0.9900.990+0.999=9901989=110221Hence,  the  required  probability  is  110221.

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
New answer posted•a year ago
S
a year ago

This is a Long Answer Type Questions as classified in NCERT Exemplar

Sol:

Let    E1  be  the  event  of  selecting  BagI           E2  be  the  event  of  selecting  BagIIand  E3  be  the  event  that  black  ball  is  selected∴       P(E1)=26=13  and  P(E2)=1−13=23  P(E3/E1)=37  and  P(E3/E2)=47∴         P(E3)=P(E1).P(E3/E1)+P(E2).P(E3/E2)                         =13.37+23.47=3+821=1121Hence,  the  required  probability  is  1121.

Thumbs Up IconUpvote Thumbs Down Icon